OpenStudy (kainui):

Let's try to prove the Goldbach Conjecture! All it says is every even number greater than 2 is the sum of 2 primes.

OpenStudy (kainui):

@M4thM1nd Alright let's do this lol.

OpenStudy (anonymous):

Here we go...

OpenStudy (kainui):

Anyone else is free to join in, we're just having fun brainstorming here. =P

OpenStudy (anonymous):

As I said, for small numbers it's easy to test all the possible combinations of two prime sums, but as these numbers get bigger... it takes too much time

OpenStudy (anonymous):

Let's try to get some insights from it

OpenStudy (kainui):

My thoughts are this: \[\Large p+q=2n \\ \Large \frac{p+q}{2}=n\] Now I can rephrase the GC as saying every number greater than 1 is the average of two prime numbers. Maybe this helps since I kind of know stuff about averages. That means that for every number n there is some point between 1 and n that reflects over to another prime:  This is basically an idea I had a while back maybe fun to think about this.